Suppose Ichiro threw a rock upward with an initial velocity of 8.0 m/s. Determine the velocity and displacement of the rock at 4 s.
1
Expert's answer
2020-02-03T14:16:49-0500

"v_0=8.0\; m/s"

"a=-9.8\; m/s^2"

"t=4\;s"

"v, \Delta y \;-\;?"

Solution:


"v=v_0+at"

"v=8.0+(-9.8)\times4=-31.2 \; m/s"


"\Delta y=v_0t+\frac{1}2at^2"

"\Delta y=8.0\times 4+\frac{1}{2}(-9.8)\times 4^2=-46.4 \;m"


Answers: "v=-31.2\; m/s, \; \Delta y=-46.4\; m."